OpenStudy (anonymous):

Suppose you draw a card from a deck of 52 cards. What is the probability that the card you draw is NOT an ace? There are 4 aces in the deck.

OpenStudy (anonymous):

Well, if you take the aces out, how many cards are left?

OpenStudy (anonymous):

48

OpenStudy (campbell_st):

P(Ace) = 4/52 so the probability of not an ace ( called the compliment) P(not Ace) = 1 - P(ace) = 1 - 4/52

OpenStudy (anonymous):

so then the chances of not picking an ace is 48 out of 52 total cards. now you can simplify that fraction.